Solve for b1: A=1/2h(b1+b2)
Answer by jim_thompson5910(35256)

A=1/2h(b1+b2)


2A=h(b1+b2)


2A=hb1+hb2


2A-hb2=hb1


(2A-hb2)/h=b1


b1 = (2A-hb2)/h


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim